OpenStudy (anonymous):

What is the reciprocal of 2 3/4?

OpenStudy (anonymous):

\[\frac{ 4 }{ 11 }\]

OpenStudy (anonymous):

Can you explain to me what you do please :)

OpenStudy (anonymous):

\[2\frac{ 3 }{ 4 }= \frac{ (4)(2)+3 }{ 4} = \frac{ 11 }{ 4}\]

OpenStudy (anonymous):

Thank you so much that really helped

OpenStudy (anonymous):

To get the inverse, literally invert the fraction

OpenStudy (anonymous):

Okay! I get it now that makes sense